OpenStudy (anonymous):

I Need Help ! Simplify completely: 5x^3+10x^2+15/5x a.x^2 + 10x^2 + 15x b.2x + 6 c.x^2 + 2x + 3 d.x^4 + 2x^3 + 3x^2

OpenStudy (cwrw238):

hint :5x^3 / 5x = x^2 so the answer begins with x^2 - now u need to divide + 10x^2 by 5x

OpenStudy (anonymous):

Okay so i did 10x^2 and i got 100 . then im suppose to divide that by 5x?

OpenStudy (cwrw238):

no pink: 10^2 = 100 this is 10 times x^2